# Hilbert-Kunz multiplicity of fiber products and idealization rings

### v3

@@ 3. Hilbert-Kunz Multiplicity of Fiber Product Rings
Let $R$, $S$, and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5) with $d:=\dim(R)\geq 2$ and $\dim R\geq\dim S>\dim T$. Suppose that $R$ is a formally unmixed local ring that is not regular. Then
$e_{HK}(R\times_{T}S)\geq 1+\frac{1}{d!(d(d+1)+1)^{d}}.$
The inequality follows easily from Proposition 3.2 and [1, Theorem 4.12].
Amalgamated duplication
Let $R$ be a Noetherian local ring. If we assume $\pi_{R}=\pi_{S}:R\to R/I$ (Diagram (2.5)) are the canonical maps and $I$ is a proper $R$-ideal, there is an isomorphism $R\times_{R/I}R\cong R\bowtie I$, where $R\bowtie I$ is the amalgamated duplication of $R$ along $I$, introduced by D’Anna [9]. Remark that, if $R$ is local with maximal ideal $\mathfrak{m}$, then $R\bowtie I$ is a local ring with maximal ideal denoted by $\mathfrak{m}\bowtie I.$ Some structural properties of $R\bowtie I$ are summarized in [2, Theorem 1.8 and 1.9 - 1.9.4].
We can use Proposition 3.2 to obtain a formula for the Hilbert-Kunz Multiplicity of the amalgamated duplication along an ideal.
3.4 Corollary.
Let $R$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be an ideal of $R$. If $\dim R/I<\dim R$, then
$e_{HK}(R\bowtie I)=2e_{HK}(R).$
Proof.
It follows directly by Proposition 3.2. ∎
3.5 Remark.
The formula given in Corollary 3.4 generalizes a formula obtained in [21, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an amalgameted duplication of $R$ along an $\mathfrak{m}$-primary ideal $I$. In this case $\dim R/I=0$ and then $e_{HK}(R\bowtie I)=2e_{HK}(R)$.
The study of structural properties of fiber product rings has been addressed by several authors, as [2] and [20]. Next corollary provides some structural properties about fiber product rings.
3.6 Corollary.
Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5) and let $I$ be an ideal of $R$. Then
  • (1)
    If $\dim R=\dim S=\dim T$ and $e_{HK}(R)+e_{HK}(S)-e_{HK}(T)>1$, then $R\times_{T}S$ is not a regular local ring.
  • (2)
    If $\dim R=\dim S>\dim T$, then $R\times_{T}S$ is not a regular local ring. In particular, if $\dim R>\dim R/I$, then $R\bowtie I$ is not a regular local ring.
  • (3)
    If $\dim R>\dim S>\dim T$ and $e_{HK}(R)>1$, then $R\times_{T}S$ is not a regular local ring.
Proof.
It is well known that regular local rings have Hilbert-Kunz multiplicity equal to one. In all the above-mentioned cases, using Proposition 3.2, we obtain $e_{HK}(R\times_{T}S)>1$ and then $R\times_{T}S$ can not be regular. ∎
As we showed above, in general, fiber product rings are not regular. The next result provides an equivalent condition for some fiber product rings to be regular.
3.7 Proposition.
Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5). If one of the following conditions holds:
  • (1)
    $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings.
  • (2)
    $R\times_{T}S$ is unmixed.
Then $R\times_{T}S$ is regular if and only if $e_{HK}(R)+e_{HK}(S)-e_{HK}(T)=1.$
Proof.
  • (1)
    By Proposition 3.2, we have $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)$. Also, we observe that as $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings, then the same is true for $R\times_{T}S$, by [2, Proposition 1.7]. For two-dimensional Cohen-Macaulay local rings, regularity is equivalent to unitary Hilbert-Kunz multiplicity, by [24, Theorem 2.5]). Then $R\times_{T}S$ is regular if and only if
    $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)=1.$
  • (2)
    The proof follows by Proposition 3.2 and [23, Theorem 1.5].
3.8 Corollary.
Let $R$, $S$ and $T$ be two-dimensional regular local rings of characteristic $p>0$ as in (2.5). Then $R\times_{T}S$ is regular.
To end this section, we will use the formulas given in Proposition 3.2 to calculate the Hilbert-Kunz multiplicity of some fiber product rings.
3.9 Example.
Let $k$ be a field of characteristic $p>0$. If $R=k[[X_{1},\dots,X_{n}]]$ and $S=k[[Y_{1},\dots,Y_{m}]]$, then
$R\times_{k}S\cong\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)},$
by Fact 2.6. Therefore, if $n>m$, then $\dim R>\dim S>\dim k$ and we have
(3.5) $e_{HK}\left(\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)}\right)=e_{HK}(K[[\underline{X}]])=1,$
by Proposition 3.2. Note that (3.5) provides examples of non-regular local rings with Hilbert-Kunz multiplicity $1$. Also, if $n=m$, then $\dim R=\dim S>\dim k$ and we have
$e_{HK}\left(\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i,j\leq n)}\right)=e_{HK}(K[[\underline{X}]])+e_{HK}(K[[\underline{Y}]])=2,$
by Proposition 3.2.
@@
$R\times_{k}S\cong\frac{k[[X,Y,Z,W,U,V]]}{(XW,XU,XV,YW,YU,YV,ZW,ZU,ZV,XY+Z^{n+1},W^{2}+UV+U^{n-1})},$
by Fact 2.6. In this case, $\dim R=\dim S>\dim k$ and by using the formulas given in [14, Example 3.18], we have
@@
for all positive integer $n\geq 4$, by Proposition 3.2.
@@
by [5, Corollary 4.2.7], taking $A=k[[X,Y]]/(X^{n}Y-Y^{2})$, $\mathfrak{p}=(X^{n}-Y)$ and $\mathfrak{q}=(Y)$. In this case, $\dim R=\dim S>\dim T$ and we have
@@
for all positive integer $n\geq 2$, by Proposition 3.2.
4. Hilbert-Kunz multiplicity of idealization rings
We will start by proving the following lemma.
@@ 4. Hilbert-Kunz multiplicity of idealization rings
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be an $R$-module. If $I$ is an ideal of $R$, $N$ is a submodule of $M$ and $J=I\ltimes N$ is an ideal of $R\ltimes M$, or equivalently $IM\subseteq N$, then
@@
$(r,m)^{q}=(r^{q},qr^{q-1}m)=(r^{q},p^{e}r^{q-1}m)=(r^{q},0).$
Then the generators of the ideal $J^{[q]}$ are of the form $(r^{q},0)$, where $r\in I$, and its elements have the form $\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)$ where $r_{i}\in I$, $s_{i}\in R$ and $m_{i}\in M$, for $i=1,\dots,n$. Observe that
$\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)=\sum_{i=1}^{n}(r_{i}^{q},0)(s_{i},m_{i})\in I^{[q]}(R\ltimes M).$
Then $J^{[q]}\subseteq I^{[q]}(R\ltimes M)$. On other hand, the generators of the $R$-module $I^{[q]}(R\ltimes M)$ are of the form $r(s,m)$, where $r\in I^{[q]}$ and $(s,m)\in R\ltimes M$. Writing $r=\sum_{i=1}^{n}y_{i}x_{i}^{q}$, where $y_{i}\in R$ and $x_{i}\in I$, for $i=1,\dots,n$, we get
@@
The next theorem provides as a consequence a formula for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes M$, where $M$ is a finite $R$-module.
@@
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. If one of the following conditions holds:
@@
    $I$ is an $\mathfrak{m}$-primary ideal of $R$ and $J=I\ltimes M$.
@@
    $I$ is an $\mathfrak{m}$-primary ideal of $R$, $N\subsetneq M$ is a submodule of $M$ which is $\mathfrak{m}$-primary and such that $J=I\ltimes N$ is an ideal of $R\ltimes M$ , or equivalently $IM\subseteq N$.
Then following equalities holds
@@
Note that in both cases the ideal $J$ is $\mathfrak{m}\ltimes M$-primary, by Theorem, 2.10. Fix $q:=p^{e}$. Using the canonical short exact sequence of $R$-modules
we get
$e_{HK}(I,R\ltimes M)=e_{HK}(I,R)+e_{HK}(I,M).$
Observe that $\ell_{R}(P)=\ell_{R\ltimes M}(P)$, for each $R\ltimes M$-module $P$. Using Lemma 4.1 we have $J^{[q]}=I^{[q]}(R\ltimes M)$ and then
@@
Also, by Theorem 2.9, we obtain the isomorphism of $R$-algebras
$\frac{R\ltimes M}{I^{[q]}\ltimes I^{[q]}M}\cong\frac{R}{I^{[q]}}\ltimes\frac{M}{I^{[q]}M}$
and an exact sequence of $R$-modules
[…]
Using length additivity in short exact sequences we obtain
@@
Now, taking limits we get the equality:
@@
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Then
@@
    If $I$ is an ideal of $R$ and $\dim R/I<\dim R$, then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
@@
    It is immediate from Theorem 2.9, by taking $J=\mathfrak{m}\ltimes R$.
@@
    Using the canonical short exact sequence of $R$-modules
    we get $e_{HK}(R)=e_{HK}(\mathfrak{m},I)+e_{HK}(\mathfrak{m},R/I)$. Since $\dim R/I<\dim R$, then $e_{HK}(\mathfrak{m},R/I)=0$ and $e_{HK}(R)=e_{HK}(\mathfrak{m},I)$. Utilizing item (1) we obtain the desired equality.
@@
The formulas given in Corollary 4.3 generalize a formula obtained in [21, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$. In this case $\dim R/I=0$ and then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
4.5 Corollary.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Then
  • (1)
    If $\dim M=\dim R$ or $e_{HK}(R)>1$, then $R\ltimes M$ is not a regular local ring.
  • (2)
    If $R$ is unmixed and there exists a finite $R$-module $M$ with $e_{HK}(R\ltimes M)=1$, then $\dim M<\dim R$ and $R$ is a regular local ring.
Proof.
  • (1)
    By Corollary 4.3, we have $e_{HK}(R\ltimes M)=e_{HK}(R)+e_{HK}(\mathfrak{m},M)>1,$ as we suppose $\dim M=\dim R$ or $e_{HK}(R)>1$. Then $R\ltimes M$ can not be regular.
  • (2)
    Using Corollary 4.3, we get $e_{HK}(R)=1$ and $e_{HK}(\mathfrak{m},M)=0$. As $R$ is unmixed it must be regular, by [23, Theorem 1.5]. Also, we have $\dim M<\dim R$, by item (3) of Proposition 2.1.
@@
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $J$ be an ideal as in Theorem 2.9. Then
@@
    where $\beta_{i}^{R}(M)$ denotes the i-th Betti number of $M$. In particular, if $M$ is a free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=(r+1)e_{HK}(I,R)$.
@@
    $e_{HK}(J,R\ltimes M)=(\operatorname{rk}_{R}(M)+1)e_{HK}(I,R),$
    Also, if in addition we assume that $M$ is a torsion-free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=e_{HK}(I,M)+re_{HK}(R)$.
@@
    Considering a short exact sequence
    we get
    $e_{HK}\left(I,R^{\mu(M)}\right)=\mu(M)e_{HK}(I,R)=e_{HK}(I,K)+e_{HK}(I,M).$
    Then $\mu(M)e_{HK}(I,R)\geq e_{HK}(I,M)$. From this inequality and by Theorem 4.2, we obtain the result.
@@
    and using the additivity of Hilbert-Kunz multiplicity on short exact sequences of finite $R$-modules, we obtain
@@
    Utilizing Theorem 2.9 we obtain the desired equality.
@@
    The proof follows by Theorem 2.9 and itens (5),(6) of Proposition 2.1.
@@
$e_{HK}\left(\frac{R[X_{1},\dots,X_{r}]}{(X_{1},\dots,X_{r})^{2}}\right)=(r+1)e_{HK}(R).$
@@
If $M=R^{r}$, then $R\ltimes M$ is naturally isomorphic to $R[X_{1},\dots,X_{r}]/(X_{1},\dots,X_{r})^{2}$, by [3, Proposition 2.2]. Therefore, the equality follows by the above corollary. ∎
5. Some bounds for the Hilbert-Kunz multiplicity
There are some fundamental conjectures about Hilbert-Kunz multiplicity to be answered. The following conjecture of Watanabe and Yoshida can be found in [23].
5.1 Conjecture.
Let $R$ be a Cohen-Macaulay local ring of characteristic $p>0$. Then for any $\mathfrak{m}$-primary ideal $I$, we have

### v4

@@ 3.1. Amalgamated duplication
To compute the Hilbert–Kunz multiplicity of the fiber product $R^{(r_{1})}\times_{k}R^{(r_{2})}$, where $R^{(r_{1})}$ and $R^{(r_{2})}$ are Veronese subrings of $R=k[[x_{1},\dots,x_{d}]]$, we use Theorem 3.2 together with the explicit formula from [31, (1.1) pg. 53]. Recall that a Veronese subring of order $r$ is defined by $R^{(r)}=\bigoplus_{n\geq 0}R_{nr},$ where $R_{nr}$ denotes the homogeneous component of degree $nr$.
The Hilbert–Kunz multiplicity of the fiber product is then given by:
$e_{HK}\big{(}R^{(r_{1})}\times_{k}R^{(r_{2})}\big{)}=\frac{1}{r_{1}}\binom{d+r_{1}-1}{r_{1}-1}+\frac{1}{r_{2}}\binom{d+r_{2}-1}{r_{2}-1}$.
Geometrically, the fiber product $R^{(r_{1})}\times_{k}R^{(r_{2})}$ can be interpreted as the union of two formal affine cones over $\mathrm{Spec}(k)$, each defined by a Veronese subring of different order.
@@
$R\times_{k}S\cong\frac{k[[x,y,z,w,u,v]]}{(xw,xu,xv,yw,yu,yv,zw,zu,zv,xy+z^{n+1},w^{2}+uv^{2}+u^{n-1})},$
by Fact 2.7. In this case, $\dim R=\dim S>\dim k$ and by using the formulas given in [20, Example 3.18], we have
@@
for all positive integer $n\geq 4$, by Theorem 3.2.
@@
by [9, Corollary 4.2.7], taking $A=k[[x,y]]/(x^{n}y-y^{2})$, $\mathfrak{p}=(x^{n}-y)$ and $\mathfrak{q}=(y)$. In this case, $\dim R=\dim S>\dim T$ and we have
@@
for all positive integer $n\geq 2$, by Theorem 3.2.
4. Hilbert-Kunz multiplicity of Nagata idealizations
We will start by proving the following lemma, which will play a crucial role in proving our main theorem of this section.
@@ 4. Hilbert-Kunz multiplicity of Nagata idealizations
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be an $R$-module. If $I$ is an ideal of $R$, NN is a submodule of $M$ and $J=I\ltimes N$ is an ideal of $R\ltimes M$, or equivalently $IM\subseteq N$, then
@@
$(r,m)^{q}=(r^{q},qr^{q-1}m)=(r^{q},p^{e}r^{q-1}m)=(r^{q},0),$
since $R$ has characteristic $p$. Then, the generators of the ideal $J^{[q]}$ are of the form $(r^{q},0)$, where $r\in I$, and its elements have the form $\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)$ where $r_{i}\in I$, $s_{i}\in R$ and $m_{i}\in M$, for $i=1,\dots,n$. Observe that
$\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)=\sum_{i=1}^{n}(r_{i}^{q},0)(s_{i},m_{i})=\sum_{i=1}^{n}r_{i}^{q}(s_{i},m_{i})\in I^{[q]}(R\ltimes M).$
Then $J^{[q]}\subseteq I^{[q]}(R\ltimes M)$. On other hand, the generators of the $R$-module $I^{[q]}(R\ltimes M)$ are of the form $r(s,m)$, where $r\in I^{[q]}$ and $(s,m)\in R\ltimes M$. Writing $r=\sum_{i=1}^{n}y_{i}x_{i}^{q}$, where $y_{i}\in R$ and $x_{i}\in I$, for $i=1,\dots,n$, we have
@@
The next theorem provides, as a consequence, a formula for the Hilbert-Kunz multiplicity of a general Nagata idealization $R\ltimes M$, where $M$ is a finitely generated $R$-module.
@@
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finitely generated $R$-module. If one of the following conditions holds:
@@
    II is an $\mathfrak{m}$-primary ideal of $R$ and $J=I\ltimes M$.
@@
    II is an $\mathfrak{m}$-primary ideal of $R$, $N\subsetneq M$ is a submodule of $M$ which is $\mathfrak{m}$-primary and such that $J=I\ltimes N$ is an ideal of $R\ltimes M$ , or equivalently $IM\subseteq N$.
The following equalities hold
@@
Note that in both cases the ideal $J$ is $\mathfrak{m}\ltimes M$-primary, by Theorem 2.12. We will fix $q:=p^{e}$. Using the canonical short exact sequence of $R$-modules
$0\rightarrow R\rightarrow R\ltimes M\rightarrow M\rightarrow 0$
we get $e_{HK}(I,R\ltimes M)=e_{HK}(I,R)+e_{HK}(I,M)$. Observe that $\ell_{R}(P)=\ell_{R\ltimes M}(P)$, for each $R\ltimes M$-module $P$ of finite lenght. Also, using Lemma 4.1 we have the equility $J^{[q]}=I^{[q]}(R\ltimes M)$ and then
@@
Also, by Theorem 2.11, we have the isomorphism
$\frac{R\ltimes M}{I^{[q]}\ltimes I^{[q]}M}\cong\frac{R}{I^{[q]}}\ltimes\frac{M}{I^{[q]}M}.$
Then, considering the exact sequence of $R$-modules of finite length
$0\rightarrow\frac{R}{I^{[q]}}\rightarrow\frac{R}{I^{[q]}}\ltimes\frac{M}{I^{[q]}M}\rightarrow\frac{M}{I^{[q]}M}\rightarrow 0$
And using length additivity in short exact sequences, we obtain
@@
Now, taking limits, we get the equality:
@@
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finitely generated $R$-module. Then
@@
    If $I$ is a proper ideal of $R$ and $\dim R/I<\dim R$, then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
@@
    It follows directly from Theorem 4.2, by taking $J=\mathfrak{m}\ltimes M$.
@@
    Using the short exact sequence $0\rightarrow I\rightarrow R\rightarrow R/I\rightarrow 0$, we get $e_{HK}(R)=e_{HK}(\mathfrak{m},I)+e_{HK}(\mathfrak{m},R/I)$. Since $\dim R/I<\dim R$, then $e_{HK}(\mathfrak{m},R/I)=0$ (see Proposition 2.1(3)) and therefore $e_{HK}(R)=e_{HK}(\mathfrak{m},I)$. Utilizing item (1), we obtain the desired equality.
@@
The formulas given in Corollary 4.3 generalize a formula obtained in [27, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$. In this case $\dim R/I=0$ and then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
@@
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finitely generated $R$-module. Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $J$ be an ideal as in Theorem 4.2. Then
@@
    where $\beta_{i}^{R}(M)$ denotes the i-th Betti number of $M$. In particular, if $M$ is a free $R$-module of rank $r$, we should have that $e_{HK}(J,R\ltimes M)=(r+1)e_{HK}(I,R)$.
@@
    $e_{HK}(J,R\ltimes M)=(\operatorname{rk}_{R}(M)+1)e_{HK}(I,R).$
@@
    Considering a short exact sequence $0\rightarrow K\rightarrow R^{\mu(M)}\rightarrow M\rightarrow 0$ we get $e_{HK}\left(I,R^{\mu(M)}\right)=\mu(M)e_{HK}(I,R)=e_{HK}(I,K)+e_{HK}(I,M)$. Then $\mu(M)e_{HK}(I,R)\geq e_{HK}(I,M)$. From this inequality and by Theorem 4.2, we obtain the result.
@@
    $0\rightarrow R^{\beta_{n}^{R}(M)}\rightarrow\cdots\rightarrow R^{\beta_{1}^{R}(M)}\rightarrow R^{\beta_{0}^{R}(M)}\rightarrow M\rightarrow 0$
    and using the additivity of Hilbert-Kunz multiplicity on short exact sequences of finitely generated $R$-modules, we obtain
@@
    Now, utilizing Theorem 4.2 we obtain the desired equality.
@@
    The proof follows by Theorem 4.2 and item (6) of Proposition 2.1.
4.6 Example.
$e_{HK}\left(\frac{R[x_{1},\dots,x_{n}]}{(x_{1},\dots,x_{n})^{2}}\right)=(n+1)e_{HK}(R).$
Indeed, if $M=R^{n}$ , then $R\ltimes M$ is naturally isomorphic to $R[x_{1},\dots,x_{n}]/(x_{1},\dots,x_{n})^{2}$, by [6, Proposition 2.2]. Therefore, the equality follows by the above corollary.
The study of structural properties of idealization rings has been addressed by several authors, as [6], [3], and [12]. The next theorem shows that an idealization ring $R\ltimes M$ is not regular. We will make use of the singularness of idealizations in the next sections.
4.7 Theorem.
Let $R$ be a Noetherian local ring and $M$ a finitely generated $R$-module. Then $R\ltimes M$ is not regular.
Proof.
It suffices to show that $R\ltimes M$ is not a domain. In fact, consider $x=(0,m)$ and $y=(0,m^{\prime}),\text{with }m,m^{\prime}\in M,\ m\neq 0,\ m^{\prime}\neq 0.$ Then $x\cdot y=(0,m)(0,m^{\prime})=(0,0\cdot m^{\prime}+0\cdot m)=(0,0).$ Therefore, $R\ltimes M$ has zero divisors and is not a domain.
@@
Let $A$ be one of the rings $R\times_{T}S$, $R\bowtie I$, or $R\ltimes M$, each of which is a three-dimensional unmixed local ring of characteristic $p>0$. Then $e_{HK}(A)\geq\frac{4}{3}$.
@@
This follows from Corollary 3.4, Theorem 4.7, and [31, Theorem 3.1 (1)]. ∎
5. Applications to the structure of singularities
In this section, we give some applications of the results obtained in the previous sections, focusing on the study of the singularities of the unmixed fiber product ring $P=R\times_{k}S$. We show, for example, that if $e_{HK}(P)=\frac{8}{3}$, then $P$ is precisely the gluing of two quadratic hypersurfaces in $\mathbb{A}^{4}_{k}$ along the origin. Moreover, we prove that if $e_{HK}(P)\in\left(\frac{8}{3},\frac{17}{6}\right],$ then $P$ is exactly the gluing of two hypersurfaces in $\mathbb{A}^{4}_{k}$ along the origin, at least one of type $x^{2}+y^{2}+z^{2}+w^{c}=0,$ for some integer $c\geq 3$. These define isolated singularities of type $A$ (or du Val singularities in dimension two), perturbed by a monomial of order at least $3$.
We also analyze the tangent structure of $P$ at its closed point. That is, we consider the associated graded ring $\operatorname{gr}_{\mathfrak{m}}(P)$, which can be interpreted as the first-order approximation of $P$ near the origin-i.e., its tangent cone.
5.1 Remark.
Let $A$ be a Noetherian local ring. Then:
[…]
Let $(A,\mathfrak{m},k)$ be any local ring of positive dimension. The associated graded ring $\operatorname{gr}_{\mathfrak{m}}(A)$ of $A$ with respect to $\mathfrak{m}$ is defined as follows:
$\operatorname{gr}_{\mathfrak{m}}(A):=\bigoplus_{n\geq 0}\mathfrak{m}^{n}/\mathfrak{m}^{n+1}.$
Then $G=\operatorname{gr}_{\mathfrak{m}}(A)$ is a homogeneous $k$-algebra such that $\mathfrak{M}=G_{+}$ is the unique homogeneous maximal ideal of $G$.
5.2 Facts.
([5, Remark 2.2]) Consider $P=R\times_{k}S$ with unique maximal ideal $\mathfrak{m}=\mathfrak{m}_{R}\oplus\mathfrak{m}_{S}$. Then